# An MLIR-based Compilation Framework for Control Flow Management on CGRAs

### v1

@@ Abstract.
Conversely, in this paper we show that control flow can be effectively managed and optimized at the compilation level, allowing for a broad set of applications to be targeted while being hardware-agnostic and achieving high performance. We embody our methodology in a modular compilation framework consisting of transformation and optimization passes, enabling support for applications with arbitrary control flows running on abstract CGRA meshes. We also introduce a novel mapping methodology that acts as a compilation back-end, addressing the limitations in available CGRA hardware resources and guaranteeing a feasible solution in the compilation process. Our framework achieves up to 2.1X speedups over state-of-the-art approaches, purely through compilation optimizations.
@@ 1. Introduction
Coarse-Grained Reconfigurable Arrays (CGRAs) are a spatial architecture composed by Processing Elements (PEs) connected in a mesh topology (839320, ; adres, ). PEs comprise ALUs and local registers, supporting arithmetic operations. Being programmable at the operation level, CGRAs present far lower overhead and much faster reconfiguration time with respect to FPGAs. Indeed, as opposed to mainstream FPGAs, CGRA typically hosts multiple configurations, one of which is selected at run-time for each clock cycle. This reconfiguration capability allows CGRAs to effectively support the acceleration of intensive parts of applications (computational kernels) defined in high-level languages (e.g., C/C++) through spatial mapping and temporal configuration. CGRAs provide a high degree of flexibility and performance, attributed to their reconfigurability and data-parallel execution model, respectively (cgra_me, ).
One of the primary challenges limiting CGRA usability is the automation of application deployment to fully exploit its parallelism. Indeed, deploying applications to effectively exploit parallelism requires substantial effort (revamp, ; 8050238, ). Most existing works simplify the CGRA compilation by limiting the compilation scope to a single DFG (chordmap, ; Plasticine, ), as highlighted in blue in Figure 1. A common approach for DFG mapping is through modulo scheduling (MS) (MeiDATE2003, ; mei2002dresc, ; SPR, ). This technique aims to overlap loop iterations, thereby minimizing the initiation intervals ($II$) - the time gap between the start of consecutive loop iterations.
Although the DFG-based mapping strategy effectively exploits the loop-level parallelism (riptide, ; compile_conclusion, ), it entirely offloads control flow management onto the host processor, as shown in Method 1 in Figure 1. Each time a kernel is deployed on the CGRA mesh, the corresponding configurations must be loaded on the hardware (10546646, ), and the input data transferred, before CGRA execution, as illustrated in Figure 2. Then, upon completion of a kernel, data must be written back into the host processor’s memory space. We herein address this shortcoming by increasing the granularity of acceleration kernels, from that of DFGs to that of entire Control Flow Graphs (CFGs).
@@
Control Data flow graph (CDFG) mapping reduces the communication overhead between basic blocks by supporting a broader scope (10.1007/978-3-030-79025-7_5, ; 7393298, ), as CDFG can span multiple basic blocks and capture both control and data dependencies. Most proposed CDFG-based approaches map each DFG within the CDFG independently (sgmf, ; 4dcgra_2019, ), as illustrated by Method 2 in Figure 1, relying on specialized hardware for managing the ensuing branches in the execution flow. When a control dependency triggers the execution of a different DFG—such as from BB0 to BB1 in Figure 1—in these works a) new instructions must be loaded for the newly scheduled blocks, and b) the relevant data must be propagated accordingly (control_plane, ). Hence, these approaches still incurs in runtime overheads and energy inefficiencies due to instruction and data reconfiguration (Snafu, ).
@@
Therefore, we argue that the limited support for complex control flows is a major limitation in state-of-the-art CGRA compilers, which are either limited in scope or adopt ad-hoc hardware solutions. Conversely, our compilation method adopts a program counter (PC)-based control flow management model. This allows the application to be compiled and configured once, eliminating the need for reconfiguration during basic block switching, as illustrated in Method 3 of Figure 1. This strategy is applicable to most CGRA architectures, which often use PC-controlled directional branches to minimize instruction memory usage (HyCUBE, ; openedge, ; morpher, ; miyamori1999remarc, ; taylor2002raw, ; revamp, ). Therein, the PC-based approach is commonly limited to the scheduling of a single loop, where it allows branching back to the entry point of the loop kernel–represented by the loop continue edge in Figure 1. We instead show that such stance can be generalized by leveraging compiler transformations, in order to accommodate arbitrary control flows.
@@
To support CDFG mapping on a PC-based controlled CGRA, a major concern is the propagation of data across basic blocks. To this end, our mapper leverages liveness analysis,  (compiler_design, ) (traditionally used in register allocation to prevent assigning the same physical register to variables that live simultaneously) to preserve data consistency on any CFGs, regardless of their sizes or connectivity. For example, in Figure 1 value ② is live-out from BB0 and live-in to BB1. Therefore, the mapping for value ③ must be placed to access value ②. Under the mapping constraints imposed by liveness analysis across CFGs, basic block transitions are executed directly via branches without incurring any reconfiguration delays, as mapping live-in/live-out values is consistently performed in-between predecessor and successor blocks.
@@
Our work considers a PC-based control model for any CDFG compilation that avoids reconfiguration overhead. Das et al. (pc_control_mapping, ) also consider a similar hardware model, in which mappings are repeatedly sampled from a search space until data consistency is achieved. Instead, we rely on liveness analysis to draw the feasible mapping space systematically. Moreover, as noted in (pc_control_mapping, ), loop-level parallelism is not supported, since a control operation must be executed at the end of each basic block to determine the subsequent execution path, thereby requiring each iteration to wait for the previous one to complete. To overcome this issue, we adopt a CFG-based transformation to adapt a new loop kernel with a loop length $L^{\prime}$ equal to $II$, as shown in Method 3 in Figure 1 Our approach ensures that our CDFG mapper provides the same scheduling efficiency of loops DFG mapper alternatives.
@@
To summarize, in this work, we present a modular compilation framework for CGRAs111 The code base is open-sourced in https://github.com/esl-epfl/Compigra.git, offering two key advantages over traditional state-of-the-art (SoA) mapping methodologies: (1) we extend the compilation scope to CDFG mapping on a PC-based hardware model, eliminating the reconfiguration overhead associated with basic block communication; and (2) we improve runtime efficiency in PC-based control model through CFG transformations to support modulo scheduling. In more detail, our main contributions are listed as follows:
@@
  • We introduce CFG optimizations applied independently at their respective levels in the compilation stack. At a higher level, basic block fusion, a hardware-agnostic technique, simplifies the CFG by merging blocks to reduce control overhead. At a lower level, modulo scheduling—a hardware-aware mapping technique—is employed to exploit loop-level parallelism.
@@ 2.1. MLIR
Multi-Level Intermediate Representation (MLIR) (MLIR, ) comprises a front-end translating high-level languages (we herein consider C/C++) into intermediate representations (IRs). In particular, MLIR supports dialects (i.e. IR variants), allowing compiler designers to define domain-specific operations, types, and attributes (ml_cgra, ; mlir_cgra, ). As a sub-project of LLVM, MLIR ensures compatibility with commonly used LLVM optimization passes and the C/C++ parsing front end. We leverage MLIR’s ability to represent both control flow and data flow at the IR level, which facilitates graph transformations in the middle end. By integrating our defined hardware-specific operations into MLIR, we achieve one-to-one operation mapping to the hardware in the back end.
@@ 2.2. Data flow graph mapping
Table 3 includes the various features supported by SoA mappers. DFG mappers, such as AA-ILP(chin_architecture-agnostic_2018, ), SAT-MapIt(satmapit, ), and E2EMap(liu_e2emap_2024, ) perform modulo scheduling, partially overlapping loop iterations, minimizing their Initiation Interval (II) to reduce runtime. Modulo scheduling handles both mapping and runtime optimization for a single loop but has limitations in managing control flow, e.g. it can only issue control-free iterations to the CGRA mesh, limiting flexibility. In these works, data transfer operations are not automated, and the DFG mappers can process only one basic block, as they do not support CFG generation, thus limiting CGRA code coverage.
Recent efforts to expand such scope include IR transformation, which merges multiple DFGs into one, employing traditional compiler optimizations such as loop unrolling and partial prediction for if-conversion (7529868, ; 10531698, ; 10.1145/2459316.2459319, ). Similarly, Dual-Issue Single-Execution (DISE) (dynamic_ii, ; compile_branch2, ) issues two branches (if-else clauses) within a single loop to the same PE, allowing the correct execution path to be determined at runtime based on the condition. While these methods extend the applicability of modulo scheduling, they do not ensure that every irregular loop can be successfully compiled—limiting the applications that can be accelerated by CGRAs (compile_branch1, ). Indeed, in these methodologies the compilation scope remains constrained, particularly for nested loops and divergent branches. Instead, we support general application compilation, enhancing mapping efficiency.
@@ 2.3. Control data flow graph management with reconfiguration
Compared to DFG mapping, the challenge of control management arises from compiling an entire CDFG composed of multiple basic blocks, each containing a DFG that captures computational dependencies, while minimizing control overhead. To expand the range of applications compiled onto CGRAs, approaches that integrate ad-hoc hardware components for control management have been proposed. In (riptide, ; cfg_intro, ), independent mapping of DFGs for each basic block is applied, while specialized hardware is incorporated for DFG switching. Dora (Dora, ) introduces dynamic binary code generation at runtime, leading to compilation and reconfiguration delays. The 4D-CGRA (4dcgra_2019, ) and Marionette PEs (control_plane, ) optimize the reconfiguration methods via specialized hardware design (see Table 3). 4D-CGRA (4dcgra_2019, ) achieves branch prediction through entry instruction matching, requiring reconfiguration for the selected block instructions and data. The marionette PE (control_plane, ) minimizes control overhead by enabling autonomous control information sharing among neighboring PEs, reducing it to a single clock cycle. However, data propagation through basic blocks remains a bottleneck for reconfiguration.
Figure 4.a illustrates the execution process for control flow management with reconfiguration. At the end of a basic block, if a control operation triggers a configuration signal from the branch operation, the control plane first reconfigures the instructions within the PEs. The instruction reconfiguration process introduces control overhead (app_stream, ; cfg_compile1, ), denoted as $D_{c}$. The required data for the newly configured basic block is either propagated from the producer PE of the previously executed block or loaded from memory, incurring additional data movement overhead, denoted as $D_{d}$. Since control flow and data flow are fully decoupled, DFG mapping is independently applied to each basic block. The innermost loop execution without control overhead is pipelined through modulo scheduling, as shown in Figures 4.c and 4.e. For sequential loops, reconfiguration is only performed during basic block switching, making it efficient with low reconfiguration overhead. However, the overhead increases significantly for nested loops. Additionally, loops with branches face a major computational bottleneck, as each basic block must be fetched separately during runtime. This limitation is illustrated in Figure 4.g.
@@ 2.4. Control data flow management using the CGRA program counter
CGRA PEs with program counter achieve basic block switching through branches  (miyamori1999remarc, ; taylor2002raw, ; revamp, ), as illustrated in Figure 4.b. Wang et al. (wang2025mlir, ) propose an integer linear programming (ILP) model for CDFG management, aiming to minimize application runtime. However, this approach suffers from exponential complexity growth as application size increases and fails to guarantee feasible solutions for general applications. Focusing on this challenge, EffiMap (pc_control_mapping, ) in Table 3 introduces a bi-directional branch in hardware to handle divergent control flows. Based on that, they introduce a heuristic scheduling method for CDFG mapping at the basic block granularity that sequentially places consumer operations to their producer PEs and stores data in internal register files. When a valid mapping cannot be found for a basic block due to register overflow or routing limitations, the algorithm backtracks to previously scheduled blocks until a feasible solution is discovered, significantly increasing compilation time. Our proposed mapping shows that CDFG can be decomposed into individual DFG mappings via liveness, as discussed in Section 7.1. Furthermore, we theoretically prove that transformations to overcome register overflow and routing limitations do not require backtracking to change established scheduling, as discussed in Section 7.4. These two strategies save compilation efficiency by decomposing the mapping complexity and avoiding rescheduling.
Moreover, their strategy results in inefficient resource utilization without exploiting loop-level parallelism. Figure 4.h illustrates this execution model for a loop with branching, where the PC dynamically controls transitions between blocks based on runtime conditions. However, since basic block execution cannot overlap under PC-based control, existing methods suffer from low scheduling efficiency. Figure 4.d and 4.f illustrate the scheduling of loops, where iterations are executed sequentially, failing to exploit loop parallelism with low hardware utilization ratio.
@@ 3. Motivation
Challenge 1: The compiler must handle diverse applications with low hardware complexity for control flow management. A general-purpose compiler should support CDFG mapping to accommodate general control flows instead of being restricted to a single DFG. State-of-the-art approaches that implement this requirement do so with dedicated CGRA control modules that handle reconfiguration and data propagation across BBs, which incur timing penalties and area cost  (control_plane, ). Compiler-based approaches can dramatically reduce both overheads, as they rely on simple program counter manipulations at run-time to handle branches (pc_control1, ). However, such a strategy has only been considered for a simple control structure, such as branching to the entry point of a control-free loop. We show that, by applying liveness analysis to basic blocks to track values propagated across them, compiler-based compilation can handle complex CDFGs with nested loops and divergent branches.
@@
Figure 5 presents examples of mapping CDFGs of applications onto a 2x2 PC-based CGRA. In particular, Figure 5.a shows a code snippet with divergent branches and its corresponding CDFG. BB0 handles the divergent branch control behavior according to the loaded value $x$, while Figure 5.b shows a mapping solution based on liveness analysis. The loaded value $x$ is the live-out of BB0 and live-in for BB1 and BB2. To ensure correct execution, the consumer operations in BB1 and BB2 must be placed on PEs that can access the live-in value $x$—specifically, PE1 for BB1 and PE0 for BB2 in the example case. Additionally, the final results of these computations, used to produce the value of $y$, must be mapped to the same hardware resources to maintain consistency. This design eliminates the need for explicitly selecting $y$’s source, as represented by the $\phi$ function in BB3. BB3 always stores the live-in value from the exact physical location (PE0 in the example), regardless of the execution path. Figure 5.d further shows the mapping solution for a nested loop shown in Figure 5.c, where the branch operations determine the control direction during runtime. The examples demonstrate that CDFG mapping can be decoupled into DFG-level compilation, provided inter-block dependencies are resolved using liveness analysis. With CFG constraints respected, each DFG can be compiled individually, as intra-block values do not affect the scheduling of other blocks.
Challenge 2: The compiler should optimize for runtime efficiency. To this end, we rely on BB fusion to increase performance, since bigger BBs offer more opportunities to exploit instruction-level parallelism. Figure 4.p illustrates an example of basic block fusion, as discussed in Section 5.1. When entire loop bodies are fused in a single BBs, loop-level parallelism can be exploited to further improve runtime efficiency. We fulfill this through CFG adaptation introduced in Section 6. Figures 4.i and 4.m illustrate such optimizations for pipelinable loops, where the adapted loop length $L$ equals the initiation interval ($II$), achieving a speedup of $L$/$II$ compared to sequential execution The optimized execution models shown in Figure 4.(j, n, and q) demonstrate higher runtime efficiency compared to both the models in Figure 4.(c, e, g) and those in Figure 4.(d, f, h).

### v2

# An MLIR-Based Compilation Framework for Control Flow Management on Coarse Grained Reconfigurable Arrays

@@ Abstract.
[…]
@@ 1. Introduction
Coarse-Grained Reconfigurable Arrays (CGRAs) are a spatial architecture composed of Processing Elements (PEs) connected in a mesh topology (839320, ; adres, ). PEs comprise ALUs and local registers, supporting arithmetic operations. Being programmable at the operation level, CGRAs present far lower overhead and much faster reconfiguration time with respect to FPGAs. Indeed, as opposed to mainstream FPGAs, CGRA typically hosts multiple configurations, one of which is selected at run-time for each clock cycle. This reconfiguration capability allows CGRAs to effectively support the acceleration of intensive parts of applications (computational kernels) defined in high-level languages (e.g., C/C++) through spatial mapping and temporal configuration. CGRAs provide a high degree of flexibility and performance, attributed to their reconfigurability and data-parallel execution model, respectively (cgra_me, ).
One of the primary challenges limiting the usability of CGRA is the automation of application deployment to fully exploit its parallelism. In fact, deploying applications to effectively exploit parallelism requires substantial effort (revamp, ; 8050238, ). Most existing works simplify the CGRA compilation by limiting the compilation scope to a single DFG (chordmap, ; Plasticine, ), as highlighted in blue in Figure 1. A common approach to DFG mapping is through modulo scheduling (MS) (MeiDATE2003, ; mei2002dresc, ; SPR, ). This technique aims to overlap loop iterations, thereby minimizing the initiation intervals ($II$) - the time gap between the start of consecutive loop iterations.
Although the DFG-based mapping strategy effectively exploits loop-level parallelism (riptide, ; compile_conclusion, ), it entirely offloads control flow management onto the host processor, as shown in Method 1 in Figure 1. Each time a kernel is deployed on the CGRA mesh, the corresponding configurations must be loaded onto the hardware (10546646, ), and the input data transferred, before CGRA execution, as illustrated in Figure 2. Then, upon completion of a kernel, the data must be written back into the host processor’s memory space. We herein address this shortcoming by increasing the granularity of acceleration kernels, from that of DFGs to that of entire Control Flow Graphs (CFGs).
@@
Control Data flow graph (CDFG) mapping reduces the communication overhead between basic blocks by supporting a larger scope (10.1007/978-3-030-79025-7_5, ; 7393298, ), as CDFG can span multiple basic blocks and capture control and data dependencies. Most of the CDFG-based approaches proposed map each DFG within the CDFG independently (sgmf, ; 4dcgra_2019, ), as illustrated by Method 2 in Figure 1, relying on specialized hardware to manage the ensuing branches in the execution flow. When a control dependency triggers the execution of a different DFG –such as from BB0 to BB1 in Figure 1—in these works (a) new instructions must be loaded for newly scheduled blocks and (b) the relevant data must be propagated accordingly (control_plane, ). Hence, these approaches still incur run-time overheads and energy inefficiencies due to instruction and data reconfiguration (Snafu, ).
@@
Therefore, we argue that the limited support for complex control flows is a major limitation in state-of-the-art CGRA compilers, which are either limited in scope or adopt ad-hoc hardware solutions. In contrast, our compilation method adopts a program counter (PC)-based control flow management model. This allows the application to be compiled and configured once, eliminating the need for reconfiguration during basic block switching, as illustrated in Method 3 of Figure 1. This strategy is applicable to most CGRA architectures, which often use PC-controlled directional branches to minimize instruction memory usage (HyCUBE, ; openedge, ; morpher, ; miyamori1999remarc, ; taylor2002raw, ; revamp, ). Therein, the PC-based approach is commonly limited to scheduling a single loop, where it allows for branching back to the entry point of the loop kernel–represented by the loop continuation edge in Figure 1. We instead show that such a position can be generalized by leveraging compiler transformations, in order to accommodate arbitrary control flows.
@@
To support CDFG mapping on a PC-based controlled CGRA, a major concern is the propagation of data across basic blocks. To this end, our mapper leverages liveness analysis  (compiler_design, ) (traditionally used in register allocation to prevent assigning the same physical register to variables that live simultaneously) to preserve data consistency in any CFG, regardless of their sizes or connectivity. For example, in Figure 1 value ② is live-out from BB0 and live-in to BB1. Therefore, the mapping of value ③ must be placed in the access value ②. Under the mapping constraints imposed by liveness analysis across CFGs, basic block transitions are executed directly via branches without incurring any reconfiguration delays, as mapping live-in/live-out values is consistently performed in-between predecessor and successor blocks.
@@
Our work considers a PC-based control model for any CDFG compilation that avoids reconfiguration overhead. Das et al. (pc_control_mapping, ) also consider a similar hardware model, in which mappings are repeatedly sampled from a search space until data consistency is achieved. Instead, we rely on liveness analysis to draw the feasible mapping space systematically. Moreover, as noted in (pc_control_mapping, ), loop-level parallelism is not supported, since a control operation must be executed at the end of each basic block to determine the subsequent execution path, thereby requiring each iteration to wait for the previous one to complete. To overcome this issue, we adopt a CFG-based transformation to adapt a new loop kernel with a loop length $L^{\prime}$ equal to $II$, as shown in Method 3 in Figure 1 Our approach ensures that our CDFG mapper provides the same scheduling efficiency of loop DFG mapper alternatives.
@@
In summary, in this work, we present a modular compilation framework for CGRAs [^1] The code base is open-sourced in https://github.com/esl-epfl/Compigra.git, offering two key advantages over traditional state-of-the-art (SoA) mapping methodologies: (1) we extend the scope of the compilation to the mapping of CDFG in a PC-based hardware model, eliminating the reconfiguration overhead associated with basic block communication; and (2) we improve the efficiency of runtime in the PC-based control model through CFG transformations to support modulo scheduling. In more detail, our main contributions are listed as follows:
@@
  • We introduce CFG optimizations applied independently at their respective levels in the compilation stack. At a higher level, basic block fusion, a hardware-agnostic technique, simplifies the CFG by merging blocks to reduce control overhead. At a lower level, modulo scheduling—a hardware-aware mapping technique—is used to exploit loop-level parallelism.
@@ 2.1. MLIR
Multi-Level Intermediate Representation (MLIR) (MLIR, ) comprises a front-end translating high-level languages (we herein consider C/C++) into intermediate representations (IRs). In particular, MLIR supports dialects (i.e. IR variants), allowing compiler designers to define domain-specific operations, types, and attributes (ml_cgra, ; mlir_cgra, ). As a sub-project of LLVM, MLIR ensures compatibility with commonly used LLVM optimization passes and the C/C++ parsing front end. We take advantage of the ability of MLIR to represent both the control flow and the data flow at the IR level, which facilitates graph transformations in the middle end. By integrating our defined hardware-specific operations into MLIR, we achieve a one-to-one operation mapping to the hardware in the back end.
@@ 2.2. Data flow graph mapping
Table 3 includes the various features supported by SoA mappers. DFG mappers, such as AA-ILP(chin_architecture-agnostic_2018, ), SAT-MapIt(satmapit, ), and E2EMap(liu_e2emap_2024, ) perform modulo scheduling, partially overlapping loop iterations, minimizing their Initiation Interval (II) to reduce runtime. Modulo scheduling handles both mapping and runtime optimization for a single loop but has limitations in managing control flow, e.g. it can only issue control-free iterations to the CGRA mesh, limiting flexibility. In these works, data transfer operations are not automated, and DFG mappers can process only one basic block because they do not support CFG generation, thus limiting the coverage of the CGRA code.
Recent efforts to expand such scope include the IR transformation, which merges multiple DFGs into one, employing traditional compiler optimizations such as loop unrolling and partial prediction for if conversion (7529868, ; 10531698, ; 10.1145/2459316.2459319, ). Similarly, Dual-Issue Single-Execution (DISE) (dynamic_ii, ; compile_branch2, ) issues two branches (if-else clauses) within a single loop to the same PE, allowing the correct execution path to be determined at runtime based on the condition. Although these methods extend the applicability of modulo scheduling, they do not ensure that every irregular loop can be successfully compiled—limiting the applications that can be accelerated by CGRAs (compile_branch1, ). Indeed, in these methodologies, the compilation scope remains constrained, particularly for nested loops and divergent branches. Instead, we support general application compilation, enhancing mapping efficiency.
@@ 2.3. Control data flow graph management with reconfiguration
Compared to DFG mapping, the challenge of control management arises from compiling an entire CDFG composed of multiple basic blocks, each containing a DFG that captures computational dependencies while minimizing control overhead. To expand the range of applications compiled onto CGRAs, approaches that integrate ad-hoc hardware components for control management have been proposed. In (riptide, ; cfg_intro, ), independent mapping of DFGs for each basic block is applied, while specialized hardware is incorporated for DFG switching. Dora (Dora, ) introduces dynamic binary code generation at runtime, leading to compilation and reconfiguration delays. The 4D-CGRA (4dcgra_2019, ) and Marionette PEs (control_plane, ) optimize the reconfiguration methods via specialized hardware design (see Table 3). 4D-CGRA (4dcgra_2019, ) achieves branch prediction through entry instruction matching, requiring reconfiguration for the selected block instructions and data. The marionette PE (control_plane, ) minimizes control overhead by enabling autonomous control information sharing among neighboring PEs, reducing it to a single clock cycle. However, data propagation through basic blocks remains a bottleneck for reconfiguration.
Figure 4.a illustrates the execution process for control flow management with reconfiguration. At the end of a basic block, if a control operation triggers a configuration signal from the branch operation, the control plane first reconfigures the instructions within the PEs. The instruction reconfiguration process introduces control overhead (app_stream, ; cfg_compile1, ), denoted as $D_{c}$. The data required for the newly configured basic block is propagated from the producer PE of the previously executed block or loaded from memory, incurring additional data movement overhead, denoted as $D_{d}$. Since control flow and data flow are fully decoupled, DFG mapping is applied independently to each basic block. The innermost loop execution without control overhead is pipelined through modulo scheduling, as shown in Figures 4.c and 4.e. For sequential loops, reconfiguration is performed only during basic block switching, making it efficient with low reconfiguration overhead. However, the overhead increases significantly for nested loops. Additionally, loops with branches face a major computational bottleneck, as each basic block must be fetched separately during runtime. This limitation is illustrated in Figure 4.g.
@@ 2.4. Control data flow management using the CGRA program counter
CGRA PEs with program counter achieve basic block switching through branches  (miyamori1999remarc, ; taylor2002raw, ; revamp, ), as illustrated in Figure 4.b. Wang et al. (wang2025mlir, ) propose an integer linear programming (ILP) model for CDFG management, aiming to minimize application runtime. However, this approach suffers from exponential complexity growth as the application size increases and does not guarantee feasible solutions for general applications. Focusing on this challenge, EffiMap (pc_control_mapping, ) in Table 3 introduces a bi-directional branch in hardware to handle divergent control flows. Based on that, they introduce a heuristic scheduling method for CDFG mapping at the basic block granularity that sequentially places consumer operations to their producer PEs and stores data in internal register files. When a valid mapping cannot be found for a basic block due to register overflow or routing limitations, the algorithm backtracks to previously scheduled blocks until a feasible solution is discovered, significantly increasing the compilation time. Our proposed mapping shows that CDFG can be decomposed into individual DFG mappings through liveness, as discussed in Section 7.1. Furthermore, we theoretically prove that transformations to overcome register overflow and routing limitations do not require backtracking to change established scheduling, as discussed in Section 7.4. These two strategies save compilation efficiency by decomposing the mapping complexity and avoiding rescheduling.
Moreover, their strategy results in inefficient resource utilization without exploiting loop-level parallelism. Figure 4.h illustrates this execution model for a loop with branching, where the PC dynamically controls the transitions between blocks based on runtime conditions. However, since basic block execution cannot overlap under PC-based control, existing methods suffer from low scheduling efficiency. Figures 4.d and 4.f illustrate loop scheduling, where iterations are executed sequentially, failing to exploit loop parallelism with a low hardware utilization ratio.
@@ 3. Motivation
Challenge 1: The compiler must handle diverse applications with low hardware complexity for control flow management. A general-purpose compiler should support CDFG mapping to accommodate general control flows instead of being restricted to a single DFG. State-of-the-art approaches that implement this requirement do so with dedicated CGRA control modules that handle reconfiguration and data propagation across BBs, incurring timing penalties and area cost  (control_plane, ). Compiler-based approaches can dramatically reduce both overheads, as they rely on simple program counter manipulations at run-time to handle branches (pc_control1, ). However, such a strategy has only been considered for a simple control structure, such as branching to the entry point of a control-free loop. We show that, by applying liveness analysis to basic blocks to track values propagated across them, compiler-based compilation can handle complex CDFGs with nested loops and divergent branches.
@@
Figure 5 presents examples of mapping CDFGs of applications on a 2x2 PC-based CGRA. In particular, Figure 5.a shows a code snippet with divergent branches and its corresponding CDFG. BB0 handles the divergent branch control behavior according to the loaded value $x$, while Figure 5.b shows a mapping solution based on liveness analysis. The loaded value $x$ is the live-out of BB0 and live-in for BB1 and BB2. To ensure correct execution, the consumer operations in BB1 and BB2 must be placed on PEs that can access the live-in value xx—specifically, PE1 for BB1 and PE0 for BB2 in the example case. Additionally, the final results of these computations, used to produce the value of $y$, must be mapped to the same hardware resources to maintain consistency. This design eliminates the need to explicitly select the source of $y$, as represented by the $\phi$ function in BB3. BB3 always stores the live-in value from the exact physical location (PE0 in the example), regardless of the execution path. Figure 5.d further shows the mapping solution for a nested loop shown in Figure 5.c, where the branch operations determine the control direction during runtime. The examples demonstrate that CDFG mapping can be decoupled into DFG-level compilation, provided inter-block dependencies are resolved using liveness analysis. With CFG constraints respected, each DFG can be compiled individually, as intra-block values do not affect the scheduling of other blocks.
Challenge 2: The compiler should optimize for runtime efficiency. To this end, we rely on BB fusion to increase performance, since larger BBs offer more opportunities to exploit instruction-level parallelism. Figure 4.p illustrates an example of basic block fusion, as discussed in Section 5.1. When entire loop bodies are fused in a single BBs, loop-level parallelism can be exploited to further improve runtime efficiency. We fulfill this through CFG adaptation introduced in Section 6. Figures 4.i and 4.m illustrate such optimizations for pipelinable loops, where the adapted loop length $L$ equals the initiation interval ($II$), achieving a speedup of LL/IIII compared to sequential execution. The optimized execution models shown in Figure 4.(j, n, and q) demonstrate higher runtime efficiency compared to both the models in Figure 4.(c, e, g) and those in Figure 4.(d, f, h).